OpenStudy (anonymous):

evaluate the following: integral S (e^3x)/((e^6x)+1) dx (hint: u=e^3x)

OpenStudy (mr.math):

\[u=e^{3x} \implies du=3e^{3x}dx\] \[I=\frac{1}{3}\int \frac{du}{u^2+1}=\frac{1}{3}\tan^{-1}u+c=\frac{1}{3}\tan^{-1}(e^{3x})+c.\]

myininaya (myininaya):

\[\int\limits_{}^{}\frac{e^{3x}}{e^{6x}+1} dx\] Thinking.... \[e^{3x}=\tan(\theta)\] => dang it

myininaya (myininaya):

great job math lol

OpenStudy (mr.math):

Obviously \(I\) is the integral you gave.

OpenStudy (mr.math):

@myin: My substitution was a hint from the question :-D

myininaya (myininaya):

\[\int\limits_{}^{}\frac{ \frac{1}{3} \sec^2(\theta)}{\tan^2(\theta)+1} d \theta\]

myininaya (myininaya):

i like my sub better

myininaya (myininaya):

it contains both subs

OpenStudy (anonymous):

haha that's awesome. i knew it was (1/3)arctan(e3x) + C. i was just seeing if this actually worked :)

OpenStudy (anonymous):

\[\int\limits \frac{e^{3 x}}{1+e^{6 x}} \, dx=\frac{1}{3} \text{ArcTan}\left[e^{3 x}\right]+c \]